# CURE: Concept Unlearning via Orthogonal Representation Editing in Diffusion Models

### v1

@@ Abstract
As Text-to-Image models continue to evolve, so does the risk of generating unsafe, copyrighted, or privacy-violating content. Existing safety interventions - ranging from training data curation and model fine-tuning to inference-time filtering and guidance - often suffer from incomplete concept removal, susceptibility to jail-breaking, computational inefficiency, or collateral damage to unrelated capabilities. In this paper, we introduce CURE, a training-free concept unlearning framework that operates directly in the weight space of pre-trained diffusion models, enabling fast, interpretable, and highly specific suppression of undesired concepts. At the core of our method is the Spectral Eraser, a closed-form, orthogonal projection module that identifies discriminative subspaces using Singular Value Decomposition over token embeddings associated with the concepts to forget and retain. Intuitively, the Spectral Eraser identifies and isolates features unique to the undesired concept while preserving safe attributes. This operator is then applied in a single step update to yield an edited model in which the target concept is effectively unlearned - without retraining, supervision, or iterative optimization. To balance the trade-off between filtering toxicity and preserving unrelated concepts, we further introduce an Expansion Mechanism for spectral regularization which selectively modulates singular vectors based on their relative significance to control the strength of forgetting. All the processes above are in closed-form, guaranteeing extremely efficient erasure in only $2$ seconds. Benchmarking against prior approaches, CURE achieves a more efficient and thorough removal for targeted artistic styles, objects, identities, or explicit content, with minor damage to original generation ability and demonstrates enhanced robustness against red-teaming. Code will be released.
@@ 1 Introduction
Recent Text-to-Image (T2I) diffusion models have garnered significant attention for their ability to synthesize high-quality, diverse images across a wide range of prompts [1, 2, 3, 4, 5, 6]. These capabilities arise from training on large-scale, uncurated internet datasets [7], which inadvertently expose models to undesirable concepts [8], such as copyrighted artistic styles [9, 10, 11, 12], deepfakes [13, 14] or inappropriate content [15, 16]. Such risks underscore the need for principled concept unlearning solutions to eliminate harmful, copyrighted, or offensive knowledge, enabling a safer and more responsible deployment of generative models.
@@
A natural first step in improving safety in generative models is to curate training data to exclude undesirable content [17]. However, retraining large models and re-annotating datasets to meet evolving safety standards is prohibitively expensive [18], and data filtering alone often leads to unintended consequences: removing one type of undesired content can expose other undesired content [19], introduce new biases [20], or result in incomplete removal [21], highlighting the limitations of data curation alone. To tackle these challenges, recent research has incorporated safety mechanisms into diffusion models. Safeguarding methods apply safety checkers to censor outputs [22], or steer generation away from unsafe concepts using classifier-free guidance and prompt filtering [16, 23]. However, these approaches introduce recurring computational overheads as guardrails must be applied for every new prompt at runtime, and lead to quality degradation in cases of prompts subjected to hard filtering due to distribution shift. Furthermore, they are easily circumvented in open-source settings where model code and parameters are publicly accessible [24, 25]. In response to the drawbacks mentioned above, an alternative is to unlearn undesirable concepts from T2I models by fine-tuning their parameters [26, 27, 28, 29, 30, 31]. Given an undesirable concept, these methods aim to prevent the generation of undesirable content by updating the model’s internal representations. Compared to full retraining or inference-time interventions, unlearning concepts by removing their knowledge from the model weights offers a more efficient and tamper-resistant solution, especially in open-source settings. However, most existing approaches require numerous fine-tuning iterations over large parameter subsets [27, 28, 29], leading to substantial computational cost and degradation in the model’s general generation quality. A compelling solution to address the above issues is offered by model editing frameworks [32, 33, 34, 35] which modify model weights in closed-form to enhance safety without the previous overheads of gradient-based fine-tuning. While more efficient, these methods often fail to completely suppress targeted inappropriate concepts, remaining vulnerable to adversarial prompts uncovered through red-teaming [36, 37, 38], which can trigger the regeneration of supposedly ‘forgotten’ content in the unlearned model. Thus, there is an urgent need for an efficient and reliable mechanism to ensure safe visual generation across a wide range of contexts.
@@
Motivated by the geometric relationship between embedding and weight spaces in diffusion models, this paper presents CURE, a mathematically principled and efficient framework for reliably erasing undesired concepts in a single-step update. Building upon prior efficient concept unlearning methods that apply closed-form solutions to modify diffusion model layers, CURE advances these approaches by explicitly exploiting orthogonal geometric structures derived through Singular Value Decomposition (SVD) [39, 40]. Specifically, we propose the Spectral Eraser, a closed-form spectral operator that constructs discriminative subspaces by decomposing token embeddings into orthogonal subspaces associated with concepts designated for removal ‘forget concepts’ and those intended to be preserved ‘retain concepts’. By translating the derived embedding-space projections directly into weight-space modifications in cross-attention layers, the Spectral Eraser precisely removes directions uniquely tied to harmful or undesired concepts while preserving the semantic integrity of overlapping and unrelated representations. Furthermore, to flexibly control the extent of concept removal, we integrate an singular-value expansion strategy, inspired by Tikhonov regularization from classical inverse problems [41, 42, 43]. This mechanism selectively scales singular vectors based on their normalized spectral energies, intuitively emphasizing or de-emphasizing the discriminative-ness of the derived forget and retain subspaces. Consequently, such spectral regularization achieves a mathematically interpretable and precise trade-off between effective concept removal and preservation of unrelated capabilities. All updates within CURE are performed in closed-form, ensuring exceptional computational efficiency, with robust unlearning in approximately $2$ seconds. Our contributions are:
@@
  • We present CURE, a strong, scalable and training-free concept unlearning method leveraging orthogonal projections and spectral geometry to yield a closed-form weight update operator, dubbed Spectral Eraser, for reliable and responsible visual content creation in T2I models.
@@ 3 Method
We present CURE, a training-free approach to concept removal in pretrained diffusion models that operates by constructing tunable geometry-aware projections using the proposed Spectral Eraser to precisely erase concepts by editing the model’s weight matrices. We begin by reviewing the core components of these models and their cross-attention mechanism, which underpin our method.
@@ 3.2 Concept Unlearning via Orthogonal Representation Editing
$\mathcal{P}_{f}=\mathcal{U}_{f}\Lambda_{f}\mathcal{U}_{f}^{\top},\quad\mathcal{P}_{r}=\mathcal{U}_{r}\Lambda_{r}\mathcal{U}_{r}^{\top},$ (5)
@@
Closed-Form Spectral Erasing Given the strength-tuned projection operators $\mathcal{P}_{f}$ and $\mathcal{P}_{r}$, we now construct a composite unlearning update that removes subspace contributions aligned with the forget set, with minimal influence on surrounding concepts by restoring overlap ($\mathcal{P}_{f}\mathcal{P}_{r}$) with the retain set:
@@
which acts on embeddings $\mathcal{E}$ to yield their updated version $\mathcal{E}^{\text{new}}=\mathcal{P}_{\text{unlearn}}*\mathcal{E}$. This closed-form solution ensures that we remove only the discriminative directions of $\mathcal{F}$ while preserving components shared with $\mathcal{R}$ – a necessary constraint to prevent performance degradation during concept erasure.
Absorbing the Unlearning Operator into Weight Space    Rather than applying $\mathcal{P}_{\text{unlearn}}$ dynamically during inference on a per-token basis, we embed this operator into the model’s stationary parameters by precomposing it directly onto the cross-attention weights. Given key and value projections are calculated as $k=W_{k}\mathcal{E}$ and $v=W_{v}\mathcal{E}$, we translate the unlearning update to the weight matrices as:
@@ 4 Experiments
Artist erasure   To evaluate the efficacy of style unlearning for mitigating artistic imitation and potential copyright violations, we follow prior works [34, 35] to use $20$ prompts each for five classical artists (Van Gogh, Pablo Picasso, Rembrandt, Andy Warhol, and Caravaggio) and five modern artists (Kelly McKernan, Thomas Kinkade, Tyler Edlin, Kilian Eng, and the series Ajin: DemiHuman), all previously reported to be mimicked by SD [12]. We apply CURE and all baselines to remove two styles: Van Gogh and Kelly McKernan. Evaluation uses LPIPS scores [66], reporting LPIPSe (on erased artists) and LPIPSu (on unerased artists), where a higher LPIPSe indicates stronger removal of the target style, and a lower LPIPSu reflects better preservation of unrelated artists. Following [23], we additionally use GPT-4o to classify artistic styles of the generated images. Acce shows how often the unlearned style is still predicted – lower is better. Accu measures accuracy on non-erased styles – higher is better. As seen in Tab. 1 and Fig. 4(a), CURE achieves effective target erasure with minimal impact on unintended styles as well as impressive specificity in preserving normal content of COCO-$30$[67], outperforming baselines. Further, CURE successfully thwarts black-box adversarial prompts crafted to trigger the ‘Van Gogh’ style, as shown in Fig.4(b), demonstrating strong robustness against red-teaming attacks [37] when others fail to resist. Finally, we evaluate our method’s scalability by erasing up to a $1000$ artist styles, while preserving all other styles. Fig. 6 shows that after $50$ erasures, outputs for the same prompt and seed start to differ, as measured by LPIPS, but CLIP scores remain stable — highlighting that the retention component of our proposed update maintains overall alignment despite perceptual changes. More results in the Appendix.
@@ 6 Conclusion
We presented CURE, a training-free, closed-form framework for fast, reliable concept unlearning in T2I diffusion models. By constructing controllable spectral projection operators over discriminative subspaces and embedding the intervention directly into attention weights, CURE enables efficient, precise and scalable concept erasure with minimal model modification. Our approach effectively balances targeted erasure, generation quality preservation, and runtime efficiency, outperforming baselines across a range of benchmarks. Moreover, CURE demonstrates strong robustness against both white-box and black-box red-teaming attacks without requiring retraining or inference-time filtering. We believe our CURE provides a principled and practical foundation for responsible deployment of generative models, thereby fostering the development of a safer AI community.
@@
This appendix provides additional technical and experimental details to support the main paper. Section 7 elaborates on the spectral expansion mechanism and its connection to classical regularization methods, along with visualizations illustrating the role of the expansion parameter $\alpha$ in controlling subspace selectivity, and hence erasure strength. Section 8 presents extended experimental results, including generalization to broader inappropriate content categories, evaluations on effective unlearning and interference on unrelated concepts, red-teaming comparisons, and assessment on general image generation quality after unlearning. This section also lists the licenses associated with all datasets and models used in this work.
7 Details on the Spectral Expansion Mechanism
7.1 Spectral Expansion function as an extension of Tikhonov regularizer
@@
7.2 Visualizing Impact of $\alpha$ in Spectral Expansion
@@
In Figure 8 (b), we visualize the heat map of the projection operator $\mathcal{P}_{f}=\mathcal{U}_{f}\Lambda_{f}\mathcal{U}_{f}^{\top}$, where $\mathcal{U}_{f}$ and the spectral energies are obtained via SVD of the forget prompt embeddings, and $\Lambda_{f}=\text{diag}(f(r_{i};\alpha))$. At low $\alpha$, weaker directions (low energy spectral components) are heavily suppressed, leading to low-rank projections focused on the highest-energy modes. As $\alpha$ increases, weaker directions are progressively included, and $\mathcal{P}_{f}$ approaches an orthogonal projector over the full concept span. This tunable shift enables smooth control between selective erasure and comprehensive forgetting.
Figure 8: (a) The spectral expansion function $f(r_{i};\alpha)$ transitions from singular-value energy-proportional weighting at $\alpha=1$ to near-uniform weighting as $\alpha\to\infty$, flattening sensitivity across modes and enabling broader subspace suppression. (b) Visualization of projection operators $\mathcal{P}_{f}=\mathcal{U}_{f}\Lambda_{f}\mathcal{U}_{f}^{\top}$, computed from the SVD of the forget prompt ‘cassette player’. For small $\alpha$, only only selective subspaces are erased, leading to weaker erasure. As $\alpha$ increases, additional modes are progressively included in the suppression, and $\mathcal{P}_{f}$ converges toward an orthogonal projector over the full forget subspace, indicating comprehensive erasure.
8 Detailed Experimental Details and Additional Results
8.1 Extended Inappropriate Content Removal

### v2

@@ Abstract
As Text-to-Image models continue to evolve, so does the risk of generating unsafe, copyrighted, or privacy-violating content. Existing safety interventions - ranging from training data curation and model fine-tuning to inference-time filtering and guidance - often suffer from incomplete concept removal, susceptibility to jail-breaking, computational inefficiency, or collateral damage to unrelated capabilities. In this paper, we introduce CURE, a training-free concept unlearning framework that operates directly in the weight space of pre-trained diffusion models, enabling fast, interpretable, and highly specific suppression of undesired concepts. At the core of our method is the Spectral Eraser, a closed-form, orthogonal projection module that identifies discriminative subspaces using Singular Value Decomposition over token embeddings associated with the concepts to forget and retain. Intuitively, the Spectral Eraser identifies and isolates features unique to the undesired concept while preserving safe attributes. This operator is then applied in a single step update to yield an edited model in which the target concept is effectively unlearned - without retraining, supervision, or iterative optimization. To balance the trade-off between filtering toxicity and preserving unrelated concepts, we further introduce an Expansion Mechanism for spectral regularization which selectively modulates singular vectors based on their relative significance to control the strength of forgetting. All the processes above are in closed-form, guaranteeing extremely efficient erasure in only $2$ seconds. Benchmarking against prior approaches, CURE achieves a more efficient and thorough removal for targeted artistic styles, objects, identities, or explicit content, with minor damage to original generation ability and demonstrates enhanced robustness against red-teaming. Project Page at https://sites.google.com/view/cure-unlearning/home.
@@ 1 Introduction
Recent Text-to-Image (T2I) diffusion models have garnered significant attention for their ability to synthesize high-quality, diverse images across a wide range of prompts [1, 2, 3, 4, 5, 6]. These capabilities arise from training on large-scale, uncurated internet datasets [7], which inadvertently expose models to undesirable concepts [8], such as copyrighted artistic styles [9, 10, 11, 12], deepfakes [13, 14] or inappropriate content [15, 16]. Such risks underscore the need for principled concept unlearning solutions to eliminate harmful, copyright-enforced, or offensive knowledge, enabling a safer and more responsible deployment of generative models.
@@
A natural first step in improving safety in generative models is to curate training data to exclude undesirable content [17]. However, retraining large models and re-annotating datasets to meet evolving safety standards is prohibitively expensive [18], and data filtering alone often leads to unintended consequences: removing one type of undesired content can expose other undesired content [19], introduce new biases [20], or result in incomplete removal [21], highlighting the limitations of data curation alone. To tackle these challenges, recent research has incorporated safety mechanisms into diffusion models. Safeguarding methods apply safety checkers to censor outputs [22], or steer generation away from unsafe concepts using classifier-free guidance and prompt filtering [16, 23]. However, these approaches introduce recurring computational overheads as guardrails must be applied for every new prompt at runtime, leading to quality degradation in cases of prompts subjected to hard filtering due to distribution shift. Furthermore, they are easily circumvented in open source settings where model code and parameters are publicly accessible [24, 25]. In response to the drawbacks mentioned above, an alternative is to unlearn undesirable concepts from T2I models by fine-tuning their parameters [26, 27, 28, 29, 30, 31]. Given an undesirable concept, these methods aim to prevent the generation of undesirable content by updating the model’s internal representations. Compared to full retraining or inference-time interventions, unlearning concepts by removing their knowledge from the model weights offers a more efficient and tamper-resistant solution, especially in open-source settings. However, most existing approaches require numerous fine-tuning iterations over large parameter subsets [27, 28, 29], leading to substantial computational cost and degradation in the model’s general generation quality. A compelling solution to address the above issues is offered by model editing frameworks [32, 33, 34, 35] which modify model weights in closed-form to enhance safety without the previous overheads of gradient-based fine-tuning. While more efficient, these methods often fail to completely suppress targeted inappropriate concepts, remaining vulnerable to adversarial prompts uncovered through red-teaming [36, 37, 38], which can trigger the regeneration of supposedly ‘forgotten’ content in the unlearned model. Thus, there is an urgent need for an efficient and reliable mechanism to ensure safe visual generation across a wide range of contexts.
@@
Motivated by the geometric relationship between embedding and weight spaces in diffusion models, this paper presents CURE, a mathematically principled and efficient framework for reliably erasing undesired concepts in a single-step update. Building upon prior efficient concept unlearning methods that apply closed-form solutions to modify diffusion model layers, CURE advances these approaches by explicitly exploiting orthogonal geometric structures derived through Singular Value Decomposition (SVD) [39, 40]. Specifically, we propose the Spectral Eraser, a closed-form spectral operator that constructs discriminative subspaces by decomposing token embeddings into orthogonal subspaces associated with concepts designated for removal, ‘forget concepts,’ and those intended to be preserved, ‘retain concepts’. By translating the derived embedding-space projections directly into weight-space modifications, the Spectral Eraser precisely removes directions uniquely tied to harmful or undesired concepts while preserving the semantic integrity of overlapping and unrelated representations. Furthermore, to flexibly control the extent of concept removal, we integrate a singular-value expansion strategy, inspired by Tikhonov regularization from classical inverse problems [41, 42, 43]. This mechanism selectively scales singular vectors based on their normalized spectral energies, emphasizing or de-emphasizing the discriminative-ness of the derived forget and retain subspaces. Consequently, such spectral regularization achieves a mathematically interpretable and precise trade-off between effective concept removal and preservation of unrelated capabilities. All updates within CURE are performed in closed-form, ensuring exceptional computational efficiency, with robust concept removal in approximately $2$ seconds. Our contributions are summarized as:
@@
  • We present CURE, a strong, scalable and training-free concept unlearning method leveraging orthogonal projections and spectral geometry, to yield a closed-form weight update operator, dubbed Spectral Eraser, for reliable and responsible visual content creation in T2I models.
@@ 3 Method
We present CURE, a training-free approach to concept removal in pretrained diffusion models that operates by constructing tunable geometry-aware projections using the proposed Spectral Eraser to precisely erase concepts by editing the model’s weight matrices. We begin by reviewing core components of these models and their cross-attention mechanism, which underpin our method.
@@ 3.2 Concept Unlearning via Orthogonal Representation Editing
$\mathcal{P}_{f}=U_{f}\Lambda_{f}U_{f}^{\top},\quad\mathcal{P}_{r}=U_{r}\Lambda_{r}U_{r}^{\top},$ (5)
@@
Closed-Form Spectral Erasing  Given the strength-tuned projection operators $\mathcal{P}_{f}$ and $\mathcal{P}_{r}$, we now construct a composite unlearning update that removes subspace contributions aligned with the forget set, with minimal influence on surrounding concepts by restoring overlap with the retain set:
@@
which acts on embeddings $\mathcal{E}$ to yield their updated version $\mathcal{E}^{\text{new}}=P_{\text{unlearn}}*\mathcal{E}$. This closed-form solution ensures that we remove only the discriminative directions of $\mathcal{F}$ while preserving components shared with $\mathcal{R}$ – a necessary constraint to prevent performance degradation during concept erasure.
Absorbing the Unlearning Operator into Weight Space   Rather than applying $P_{\text{unlearn}}$ dynamically during inference on a per-token basis, we embed this operator into the model’s stationary parameters by precomposing it directly onto the cross-attention weights. Given key and value projections are calculated as $k=W_{k}\mathcal{E}$ and $v=W_{v}\mathcal{E}$, we translate the unlearning update to the weight matrices as:
@@ 4 Experiments
Artist erasure  To evaluate the efficacy of style unlearning for mitigating artistic imitation and potential copyright violations, we follow prior works [34, 35] to use $20$ prompts each for five classical artists (Van Gogh, Pablo Picasso, Rembrandt, Andy Warhol, and Caravaggio) and five modern artists (Kelly McKernan, Thomas Kinkade, Tyler Edlin, Kilian Eng, and the series Ajin: DemiHuman), all previously reported to be mimicked by SD [12]. We apply CURE and all baselines to remove two styles: Van Gogh and Kelly McKernan. Evaluation uses LPIPS scores [66], reporting LPIPSe (on erased artists) and LPIPSu (on unerased artists), where a higher LPIPSe indicates stronger removal of the target style, and a lower LPIPSu reflects better preservation of unrelated artists. Following [23], we additionally use GPT-4o to classify artistic styles of the generated images. Acce shows how often the unlearned style is still predicted – lower is better. Accu measures accuracy on non-erased styles – higher is better. As seen in Tab. 1 and Fig. 4(a), CURE achieves effective target erasure with minimal impact on unintended styles as well as impressive specificity in preserving normal content of COCO-3030[67], outperforming baselines. Further, CURE successfully thwarts black-box adversarial prompts crafted to trigger the ‘Van Gogh’ style, as shown in Fig.4(b), demonstrating strong robustness against red-teaming attacks [37] when others fail to resist. Finally, we evaluate our method’s scalability by erasing up to $1000$ artist styles, while preserving all other styles. Fig. 6 shows that after $50$ erasures, outputs for the same prompt and seed start to differ, as measured by LPIPS, but CLIP scores remain stable — highlighting that the retention component of our proposed update maintains overall alignment despite perceptual changes. More results in the Appendix.
@@ 6 Conclusion
We present CURE, a training-free, closed-form framework for fast, reliable concept unlearning in T2I diffusion models. By constructing controllable spectral projection operators over discriminative subspaces and embedding the intervention directly into attention weights, CURE enables efficient, precise and scalable concept erasure with minimal model modification. Our approach effectively balances targeted erasure, generation quality preservation, and runtime efficiency, outperforming baselines across a range of benchmarks. Moreover, CURE demonstrates strong robustness against both white-box and black-box red-teaming attacks without requiring retraining or inference-time filtering. We believe CURE provides a principled and practical foundation for responsible deployment of generative models, thereby fostering the development of a safer AI community.
7 Potential Impact Statement
CURE can erase any concept that admits a sufficiently precise textual representation in the model’s embedding space. This dependency makes arbitrary removal of unknown backdoors or watermarks non-trivial: an adversary must already know what they intend to target and be able to specify it textually. However, as with other editing tools, misuse risks exist. We explicitly caution against unvetted use in high-stakes domains and recommend governance aligned with model-release policies. To this end, we outline safeguards that practitioners can layer atop CURE: (i) Prompt filtering to prevent unlearning of protected tokens or safety-critical concepts; (ii) Gated execution in deployment, where CURE is sandboxed and permitted only for a whitelisted set of concepts; (iii) Auditability via logging of edits and publishing of the exact projection operator thereof for reproducibility and oversight. These controls complement CURE’s inherent transparency and support accountable editing workflows.
8 Acknowledgment
This work was supported in part by the Center for the Co-Design of Cognitive Systems (CoCoSYS), a research center under the Joint University Microelectronics Program (JUMP) 2.0, a Semiconductor Research Corporation (SRC) initiative sponsored by DARPA; and by the U.S. Department of Energy (DOE) Office of Science, Office of Basic Energy Sciences, and by National Science Foundation.
@@ 8 Acknowledgment
This appendix provides additional technical and experimental details to support the main paper. Section A elaborates on the spectral expansion mechanism and its connection to classical regularization methods, along with visualizations illustrating the role of the expansion parameter $\alpha$ in controlling subspace selectivity, and hence erasure strength. Section B presents extended experimental results, including generalization to broader inappropriate content categories, evaluations on effective unlearning and interference on unrelated concepts, red-teaming comparisons, and assessment on general image generation quality after unlearning. Next, Sec. C provides additional discussions for CURE and experimental choices. Finally, Sec. D lists the licenses associated with all datasets and models used in this work.
Appendix A Details on the Spectral Expansion Mechanism
A.1 Spectral Expansion function as an extension of Tikhonov regularizer
@@
A.2 Visualizing Impact of $\alpha$ in Spectral Expansion
@@
In Figure 8 (b), we visualize the heat map of the projection operator $\mathcal{P}_{f}=U_{f}\Lambda_{f}U_{f}^{\top}$, where $U_{f}$ and the spectral energies are obtained via SVD of the forget prompt embeddings, and $\Lambda_{f}=\text{diag}(f(r_{i};\alpha))$. At low $\alpha$, weaker directions (low energy spectral components) are heavily suppressed, leading to low-rank projections focused on the highest-energy modes. As $\alpha$ increases, weaker directions are progressively included, and $\mathcal{P}_{f}$ approaches an orthogonal projector over the full concept span. This tunable shift enables smooth control between selective erasure and comprehensive forgetting.
A.3 Effect of the Spectral Suppression Strength $\alpha$
We study how the scalar suppression strength $\alpha$ in our Spectral Eraser controls the erase-retain trade-off. We run “cat” removal over a 100-image set spanning {tiger, lion, cheetah, leopard, dog, rabbit, mouse, bird, cat, feline}. Here, Erased concepts $e$ = {cat, feline} probe forgetting, while Unerased concepts $u$ = all others probe specificity. We report LPIPS ($\uparrow$ better for $e$, $\downarrow$ better for $u$) and a GPT-4o classifier accuracy ($\downarrow$ better for $e$, $\uparrow$ better for $u$) as a semantic erasure proxy.
$\alpha$ LPIPSe $\uparrow$ LPIPSu $\downarrow$ Acce $\downarrow$ Accu $\uparrow$
1 0.41 0.17 0.47 0.96
2 0.46 0.19 0.08 0.94
5 0.47 0.26 0.06 0.86
10 0.49 0.27 0.05 0.85
100 0.51 0.30 0.00 0.86
1000 0.58 0.31 0.00 0.85
$\infty$ 0.65 0.34 0.00 0.52
Table 6: Ablation of spectral suppression $\alpha$. Larger $\alpha$ drives stronger forgetting (low Acce) but hurts unrelated concepts (higher LPIPSu, lower Accu). A moderate value balances both.
[…]
Figure 8: (a) The spectral expansion function $f(r_{i};\alpha)$ transitions from singular-value energy-proportional weighting at $\alpha=1$ to near-uniform weighting as $\alpha\to\infty$, flattening sensitivity across modes and enabling broader subspace suppression. (b) Visualization of projection operators $\mathcal{P}_{f}=U_{f}\Lambda_{f}U_{f}^{\top}$, computed from the SVD of the forget prompt ‘cassette player’. For small $\alpha$, only only selective subspaces are erased, leading to weaker erasure. As $\alpha$ increases, additional modes are progressively included in the suppression, and $\mathcal{P}_{f}$ converges toward an orthogonal projector over the full forget subspace, indicating comprehensive erasure.
Appendix B Detailed Experimental Details and Additional Results
B.1 Extended Inappropriate Content Removal